### v1

# Experimentally Tractable Generation of High-Order Rogue Waves in Bose-Einstein Condensates

Abstract
In this work we study a prototypical, experimentally accessible scenario that enables the systematic generation of so-called high-order rogue waves in atomic Bose-Einstein condensates. These waveforms lead to significantly and controllably more extreme focusing events than the famous Peregrine soliton. In one spatial dimension, we showcase conclusive numerical evidence that our scheme generates the focusing behavior associated with the first four rogue waves from the relevant hierarchy. We then extend considerations to anisotropic two-dimensional and even three-dimensional settings, establishing that the scheme can generate second order rogue waves despite the well-known limitation of finite-time blow up of focusing nonlinear Schrödinger equations.
Introduction. Over the last two decades, the study of rogue waves has become one of the most active themes of study within nonlinear science Yan (2012); Onorato et al. (2013); Dudley et al. (2014); Mihalache (2017); Dudley et al. (2019); Tikan et al. (2021). A series of remarkable developments, initially in the field of nonlinear optics Solli et al. (2007, 2008); Kibler et al. (2010, 2012); DeVore et al. (2013); Frisquet et al. (2016); Tikan et al. (2017), led to detection tools for probing rogue waves and also suggested their relevance in other applications, such as in supercontinuum generation. Meanwhile, state-of-the-art experiments in fluid mechanics Chabchoub et al. (2011, 2012); Chabchoub and Fink (2014) not only enabled the realization of the famous prototypical Peregrine soliton nonlinear waveform Peregrine (1983), but also a higher-order breather referred to as a “super rogue wave” Chabchoub et al. (2012). Importantly, other areas of dispersive wave dynamics including plasmas Bailung et al. (2011); Sabry et al. (2012); Tolba et al. (2015) and more recently atomic Bose-Einstein condensates Romero-Ros et al. (2024) have offered additional fertile platforms for the exploration of associated rogue wave dynamics.
What is perhaps less familiar in the physics community is how these rogue waves naturally emerge in different wave-breaking scenarios of dispersive wave PDE models. Here, we argue that such an understanding proves essential to their physical implementation and immediately provides an experimentally tractable example. Starting with the universal Korteweg-de Vries equation, a famous Dubrovin conjecture proved in Claeys and Grava (2009) characterized wave breaking as described by a solution of the second member of the Painlevé-I hierarchy. Similarly, for the universal nonlinear Schrödinger (NLS) model, the proof of another Dubrovin conjecture in Bertola and Tovbis (2013) established the wave-breaking description via a solution of the the tritronquée solution of the Painlevé-I equation. The work of Lu and Miller (2022) obtained similar results for the Sine-Gordon model.
Our focus herein involves a less generic, yet still very much experimentally tractable scenario involving the NLS equation; a ubiquitous physics model that arises in nonlinear optics Kivshar and Agrawal (2003); Hasegawa and Kodama (1995), atomic physics Pitaevskii and Stringari (2003); Pethick and Smith (2002); Kevrekidis et al. (2015) and plasma dynamics Kono and Skorić (2010). The NLS equation features a non-generic focusing example based on a Painlevé-III type breaking known to arise from the so-called Talanov, or semi-circle, initial data Talanov (1970); Gurevich and Shvartsburg (1970). As recently explained in Demontis et al. (2023), the dispersionless limit of this problem exhibits a finite-time blow up with the presence of dispersion manifesting so-called high-order rogue waves (HORWs). Reference He et al. (2013) shows how HORWs can be obtained via Darboux transformations while references Bilman et al. (2020); Bilman and Miller (2022) provide a more mathematically detailed study thereof. Indeed, this process has an asymptotic limit, namely the so-called infinite-order rogue wave that was identified, for the first time to our knowledge, in the work of Suleimanov (2017).
Our aim is to show that atomic Bose-Einstein condensates present an excellent opportunity for the generation of high-order, and potentially infinite order, rogue waves dynamically. This is due to the fact that in the well-known Thomas-Fermi (TF) limit of large density, such systems under self-defocusing (self-repulsive) nonlinearity acquire a well-established parabolic density profile Pitaevskii and Stringari (2003); Pethick and Smith (2002); Kevrekidis et al. (2015), up to a small boundary-layer and curvature-driven correction that has been analyzed in Karali and Sourdis (2015); Gallo and Pelinovsky (2009).
This provides approximately semi-circular initial wavefunction data “for free”, as it is the ground state of the self-repulsive problem provided that one can perform a quench in the nonlinearity from the defocusing to the focusing, modulationally unstable regime. Fortunately, such quenches are known to be quite feasible in atomic BECs, from the early works of Cornish et al. (2000); Roberts et al. (2001), demonstrations of wide tunability in Pollack et al. (2009), and even the recent realization of Townes soliton collapse in Chen and Hung (2020). Indeed, the very recent work of Banerjee et al. (2024) has used such a quench to probe the collapse of a vortical BEC pattern.
We show that this technique can be tuned through the chemical potential, the trap strength, and the scattering length to provide good approximations of $k^{\rm th}$-order rogue waves for $k=2$, $k=3$ and $k=4$, with the potential to, in principle, generalize this approach to higher orders. We then go beyond one-dimensional settings, as is more realistic in atomic BECs, and illustrate that even higher-dimensional settings with a quasi-1D confinement still provide the possibility of an excitation of higher order rogue waves such as $k=2$ and $k=3$. This, in turn, strongly suggests that many of the above experiments bear the controllable formation of such higher-order rogue waves at their fingertips and the experimental range to which this technology can be used to obtain excitations of large amplitude waveforms still remains to be (hopefully, imminently) explored.
Theoretical Analysis of HORWs. To briefly provide the theoretical background of the type of rogue waves that we consider throughout this work, we closely follow the concise account of Bilman, et al. Bilman et al. (2020). We begin by considering the prototypical focusing NLS equation
$i{\partial_{t}\psi}=-\frac{1}{2}{\partial_{x}^{2}\psi}-|\psi|^{2}\psi,$ (1)
where $x\in\mathbb{R}.$ A potential term $V(x)\psi$ added to the right hand side of (1) will also (when appropriate) be considered in the numerical examples that follow. Notice that for all considerations below, for breadth of exposition, we will maintain our quantities dimensionless. Numerous systematic topical expositions, e.g., in atomic physics Pethick and Smith (2002); Pitaevskii and Stringari (2003); Kevrekidis et al. (2015) or optics Kivshar and Agrawal (2003); Hasegawa and Kodama (1995) detail how to translate these general findings to dimensional units, as needed. An expression of the HORWs that exactly solve the NLS equation (1) can be constructed through the following procedure.
Consider the expansion coefficients from the suitable series expansion around $\lambda=i$, $F_{\ell}(x,t)$ and $G_{\ell}(x,t),\mathrm{where}\ \ell\in\mathbb{Z}^{+}$, arising from the generating functions
$(1-i\lambda)\frac{\sin\left((x+\lambda t)\sqrt{\lambda^{2}+1}\right)}{\sqrt{\lambda^{2}+1}}$ $=\sum_{\ell=0}^{\infty}\left(\frac{i}{2}\right)^{\ell}F_{\ell}(x,t)(\lambda-i)^{\ell},$
$\cos\left((x+\lambda t)\sqrt{\lambda^{2}+1}\right)$ $=\sum_{\ell=0}^{\infty}\left(\frac{i}{2}\right)^{\ell}G_{\ell}(x,t)(\lambda-i)^{\ell}.$
We use these coefficients to define the following $k\times k$ matrices
$K_{pq}^{(k)}(x,t):=\sum_{\mu=0}^{p-1}\sum_{v=0}^{q-1}\left(\begin{array}[]{c}\mu+v\\ \mu\end{array}\right)\left(F^{*}_{q-v-1}F_{p-\mu-1}+G^{*}_{q-v-1}G_{p-\mu-1}\right),$
and
$H_{pq}^{(k)}(x,t):=-2\left(F_{p-1}+G_{p-1}\right)\left(F^{*}_{q-1}-G^{*}_{q-1}\right),$
where $1\leq p,q\leq k$ and with * denoting complex conjugation, while the arguments of $F$ and $G,$ being $(x,t),$ are implied for brevity. Rogue waves of order $k$, denoted by $\psi_{k}(x,t),$ are thus furnished by the following formula
$\psi_{k}(x,t):=(-1)^{k}\frac{\operatorname{det}\left(\mathbf{K}^{(k)}(x,t)+\mathbf{H}^{(k)}(x,t)\right)}{\operatorname{det}\left(\mathbf{K}^{(k)}(x,t)\right)}.$ (2)
The compact representation afforded by Equation (2) coincides with the leading order solution of a parametrized matrix Riemann-Hilbert problem. See  Bilman et al. (2020), in particular Proposition 1, for more details.
The $k=1$ Peregrine solution, up to an unimportant phase factor, is given by
$\psi_{1}(x,t):=1-4\frac{1+2\mathrm{i}t}{1+4x^{2}+4t^{2}}.$
We show the space-time evolution of the first four of these rogue waves, in absolute value, in Figure 1. Note that this process can be generalized for arbitrary $k$, leading eventually to the limit of $k\rightarrow\infty$ which corresponds to the waveform discovered in Suleimanov (2017). The sequence of the states, through having $k$ “stems” and from which an extreme focusing emerges, is transparent in the space-time evolution of the patterns in Figure 1.
Figure 1: Visualization of the NLS rogue waves, in absolute value, whose formulae are given by Equation (2) and the immediately preceding definitions. The top panel shows the rogue waves’ spatial profiles at $t=0$ coinciding with their peak focusing time. The bottom one shows their space-time contour plot evolution.
Numerical Examples: the One-Dimensional Case. We now turn to the one-dimensional realization of our proposed experimental protocol (within atomic BECs Pethick and Smith (2002); Pitaevskii and Stringari (2003); Kevrekidis et al. (2015), or, by extension, optical media with parabolic refractive index; see, e.g., Longhi (2003); Conforti et al. (2017) for some relevant examples). We start at the defocusing regime with a parabolic confinement $V(x)=(1/2)\Omega^{2}x^{2}$ and nonlinear defocusing prefactor $g_{\rm defocus}$, for a controllable chemical potential $\mu$; both $\mu$ and $\Omega$ are tunable, and $\mu\gg\Omega$ defines the effective TF regime. The ground state wavefunction is approximately semicircular with density controllably approximated by $|\psi_{TF}|^{2}=\mu-V(x)$, where $\mu>V(x)$. The approximation becomes progressively better as the TF regime is approached, with an accordingly shrinking boundary layer (see, also Karali and Sourdis (2015); Gallo and Pelinovsky (2009)). Now, at $t=0$, we switch the nonlinearity to a focusing one of prefactor $g_{\rm focus}$, i.e., solving Eq. (1), but with (again, controllable attractive interactions Donley et al. (2001); Roberts et al. (2001); Pollack et al. (2009)) nonlinear prefactor $g_{\rm focus}$ and with this controllably close to a semi-circular initial condition.
To make the claim that this protocol enables the controlled generation (as the parameters $\mu$, $g$ and $\Omega$ are varied) of high-order rogue waves, we perform a fit of our numerical realizations thereof, at the time of maximal focusing, to the exact solutions given by Equation (2) rescaled appropriately in space and by the value of the nonlinear focusing coefficient $g_{\rm focus}$ such that the NLS equation (1) remains invariant.
We denote the rescaling in space used to minimize the mismatch in the ${L^{2}(\mathbb{R})}$ norm between the numerics and the exact solution by the symbol $s.$ Since the functions given by Equation (2) have a non-zero background, we minimize the mismatch centered at space and up until the absolute value reaches its last local minimum to the left and to the right of the origin. In the figure captions, we report the parameters $\left(\Omega,g_{\rm defocus},g_{\rm focus},\mu,s\right)$ found that minimize the mismatch between numerical simulations and the exact formula given by Equation (2).
In Figure 2, we show the match between the first four generated rogue waves and the exact expressions from Equation (2). Naturally, and similarly to what occurs in the case of the single Peregrine in BECs Romero-Ros et al. (2024), we can only hope to locally match the $k$-th order rogue waves, given the distinct nature of the spatial asymptotics of our initial conditions, hence the need for our proposed matching selection. Nevertheless, our fitting procedure clearly suggests the local spontaneous generation of the relevant multi-hump patterns, by analogy with what happens for Peregrine-based wave breaking in Bertola and Tovbis (2013).
Figure 2: Local fitting of our numerical simulations to the first $4$ (i.e., $k=1$ to $k=4$) HORWs given by Equation (2). We report the parameters $\left(\Omega,g_{\rm defocus},g_{\rm focus},\mu,s\right)$ to three digits of precision. Top left: $(1.06,0.87,-0.443,3.42,0.914)$. Top right: $(1.23,0.891,-0.701,6.19,1.74)$. Bottom Left: $(1.25,1.07,-0.794,7.42,1.49)$. Bottom Right: $(1.93,1.84,-.483,18.2,1.01)$.
Higher-dimensional Generalizations. Extending considerations to two spatial dimensions, we are able to observe HORWs, e.g., with $k=2$ or $k=3$ in anisotropic variants of the latter setting. More concretely, we show the formation of focusing strongly reminiscent of a second order rogue wave in the top panel of Figure 3. It is clear here that we need an experimentally accessible, highly anisotropic 2D trap $V(x,y)=\frac{\Omega_{x}^{2}}{2}x^{2}+\frac{\Omega_{y}^{2}}{2}y^{2}$ in order to achieve the relevant dynamics. We further show the shape of the initial condition used to illustrate the level of anisotropy used to ensure the prevention of finite-time blowup of the dynamics. During the focusing dynamics, we release the trap in the $x-$direction while maintaining the trap in the transverse, that is, $y-$direction fixed. In the bottom panel of Figure 3, we show the formation of a $k=3$ rogue wave, as well as an isosurface of the two-dimensional dynamics.
[…]
In three spatial dimensions, anisotropy can once again serve to mitigate the risk of collapse, as is well-known, e.g., from works such as Pérez-García et al. (1997); Salasnich et al. (2002); Márquez and Evaristo (2014). This, in turn, enables us, within the regimes of sufficiently tightly transversely confined BECs discussed therein, to identify parameters that lead to a genuine second order rogue wave-type scenario. This is shown in Fig. 4 both at the level of the quasi-1D space-time evolution (left panel) and at that of the detailed comparison of the main feature in a quasi-1D integrated form with the 2nd order rogue wave of the Equation (2).
Conclusions and Future Challenges. In the present work we revisited the framework of higher-order rogue waves that has been recently developed in the mathematical literature (via Darboux transformations and robust inverse scattering methods) as a non-generic, yet well-accessible form of wave breaking. While, to the best of our knowledge, this setup has not been explored at the level of realistic or experimentally relevant settings, we have showcased examples that are fully accessible to current atomic BEC experiments (and also elsewhere, such as, e.g., suitably designed optical media). We then went on to explore the freedom available to the system through a variation of parameters to provide scenarios enabling the realization of $k=1$ through $k=4$ rogue waves, with a view towards higher order rogue waves, including the intriguing infinite-order one of Suleimanov (2017). While the relevant idea is, in its essence, tailored to one spatial dimension, we have illustrated that higher-dimensional, highly anisotropic settings are also very much accessible and conducive towards such considerations.
Naturally, this study paves the way for numerous further considerations of related concepts. We only mention a few of these here. One necessity is a wide parametric mapping of the space of possible outcomes of near-semi-circular initial data in the space of $(\mu,\Omega)$ etc. Another question: how amenable are non-integrable systems to such types of wave breaking? For example, will HORW-type behavior survive in quintic, cubic-quintic, or other recently proposed so-called droplet Luo et al. (2021) variants of the model? A further direction stems from spatially discrete models and concerns whether a variant of such higher-order (including infinite-order Suleimanov (2017)) rogue waves exists of Ablowitz-Ladik type and discrete nonlinear Schrödinger type lattices Ablowitz et al. (2004); Kevrekidis (2009).
It can thus be inferred that this theme presents a wide palette of opportunities in one- and higher- spatial dimensions that are extremely timely and relevant for near-future analytical, numerical and experimental considerations.