### v3

# Experimentally Tractable Generation of High-Order Rogue Waves in Bose-Einstein Condensates

Abstract
In this work, we study a prototypical, experimentally accessible scenario that enables the systematic generation of so-called high-order rogue waves in atomic Bose-Einstein condensates. These waveforms lead to significantly and controllably more extreme focusing events than the famous Peregrine soliton.
In one spatial dimension, we showcase conclusive numerical evidence that our scheme generates the focusing behavior associated with the first four rogue waves from the relevant hierarchy. We then extend considerations to anisotropic two-dimensional and even three-dimensional settings, establishing that the scheme can generate second order rogue waves despite the well-known limitation of finite-time blow up of focusing nonlinear Schrödinger equations.
Introduction. Over the last two decades, the study of rogue waves has become one of the most active themes of study within nonlinear science Yan (2012); Onorato et al. (2013); Dudley et al. (2014); Mihalache (2017); Dudley et al. (2019); Tikan et al. (2021). A series of remarkable developments, initially in the field of nonlinear optics Solli et al. (2007, 2008); Kibler et al. (2010, 2012); DeVore et al. (2013); Frisquet et al. (2016); Tikan et al. (2017), led to detection tools for probing rogue waves and also suggested their relevance in other applications, such as in supercontinuum generation. Meanwhile, state-of-the-art experiments in fluid mechanics Chabchoub et al. (2011, 2012); Chabchoub and Fink (2014) not only enabled the realization of the famous prototypical Peregrine soliton nonlinear waveform Peregrine (1983), but also a higher-order breather referred to as a “super rogue wave” Chabchoub et al. (2012). Importantly, other areas of dispersive wave dynamics including plasmas Bailung et al. (2011); Sabry et al. (2012); Tolba et al. (2015) and more recently atomic Bose-Einstein condensates Romero-Ros et al. (2024) have offered additional fertile platforms for the exploration of associated rogue wave dynamics.
What is perhaps less familiar in the physics community is how these rogue waves naturally emerge in different wave-breaking scenarios of dispersive wave PDE models. Here, we argue that such an understanding proves essential to their physical implementation and immediately provides an experimentally tractable example. Starting with the universal Korteweg-de Vries equation, a famous Dubrovin conjecture proved in Claeys and Grava (2009) characterized wave breaking as described by a solution of the second member of the Painlevé-I hierarchy. Similarly, for the universal nonlinear Schrödinger (NLS) model, the proof of another Dubrovin conjecture in Bertola and Tovbis (2013) established the wave-breaking description via a solution of the the tritronquée solution of the Painlevé-I equation. The work of Lu and Miller (2022) obtained similar results for the Sine-Gordon model.
Our focus herein involves a less generic, yet still very much experimentally tractable scenario involving the NLS equation; a ubiquitous physics model that arises in nonlinear optics Kivshar and Agrawal (2003); Hasegawa and Kodama (1995), atomic physics Pitaevskii and Stringari (2003); Pethick and Smith (2002); Kevrekidis et al. (2015) and plasma dynamics Kono and Skorić (2010). The NLS equation features a non-generic focusing example based on a Painlevé-III type breaking known to arise from the so-called Talanov, or semi-circle, initial data Talanov (1970); Gurevich and Shvartsburg (1970). As recently explained in Demontis et al. (2023), the dispersionless limit of this problem exhibits a finite-time blow up with the presence of dispersion manifesting so-called high-order rogue waves (HORWs). Reference He et al. (2013) shows how HORWs can be obtained via Darboux transformations while references Bilman et al. (2020); Bilman and Miller (2022) provide a more mathematically detailed study thereof; see also Akhmediev (2021) for a more physically minded review. Indeed, this process has an asymptotic limit, namely the so-called infinite-order rogue wave that was identified, for the first time to our knowledge, in the work of Suleimanov (2017).
Our aim is to show that atomic Bose-Einstein condensates present an excellent opportunity for the generation of high-order, and potentially infinite order, rogue waves dynamically. This is due to the fact that in the well-known Thomas-Fermi (TF) limit of large density, such systems under self-defocusing (self-repulsive) nonlinearity acquire a well-established parabolic density profile Pitaevskii and Stringari (2003); Pethick and Smith (2002); Kevrekidis et al. (2015), up to a small boundary-layer and curvature-driven correction that has been analyzed in Karali and Sourdis (2015); Gallo and Pelinovsky (2009).
This provides approximately semi-circular initial wavefunction data “for free”, as it is the ground state of the self-repulsive problem provided that one can perform a quench in the nonlinearity from the defocusing to the focusing, modulationally unstable regime. Fortunately, such quenches are known to be quite feasible in atomic BECs, from the early works of Cornish et al. (2000); Roberts et al. (2001), demonstrations of wide tunability in Pollack et al. (2009), and even the recent realization of Townes soliton collapse in Chen and Hung (2020). Indeed, such a quench was proposed towards creating a fundamental (Peregrine) rogue wave in Bludov et al. (2009). Furthermore, the very recent work of Banerjee et al. (2024) has used such a quench to probe the collapse of a vortical BEC pattern.
We show that this technique can be tuned through the chemical potential, the trap strength, and the scattering length to provide good approximations of $k^{\rm th}$-order rogue waves for $k=2$, $k=3$ and $k=4$, with the potential to, in principle, generalize this approach to higher orders. We then go beyond one-dimensional settings, as is more realistic in atomic BECs, and illustrate that even higher-dimensional settings with a quasi-1D confinement still provide the possibility of an excitation of higher order rogue waves such as $k=2$ and $k=3$. This, in turn, strongly suggests that many of the above experiments bear the controllable formation of such higher-order rogue waves at their fingertips and the experimental range to which this technology can be used to obtain excitations of large amplitude waveforms still remains to be (hopefully, imminently) explored.
Theoretical Analysis of HORWs. To briefly provide the theoretical background of the type of rogue waves that we consider throughout this work, we closely follow the concise account of Bilman, et al. Bilman et al. (2020). We begin by considering the prototypical focusing NLS equation
$i{\partial_{t}\psi}=-\frac{1}{2}{\partial_{x}^{2}\psi}-|\psi|^{2}\psi,$ (1)
where $x\in\mathbb{R}.$ A potential term $V(x)\psi$ added to the right hand side of (1) will also (when appropriate) be considered in the numerical examples that follow. Notice that for all considerations below, for breadth of exposition, we will maintain our quantities dimensionless. Numerous systematic topical expositions, e.g., in atomic physics Pethick and Smith (2002); Pitaevskii and Stringari (2003); Kevrekidis et al. (2015) or optics Kivshar and Agrawal (2003); Hasegawa and Kodama (1995) detail how to translate these general findings to dimensional units, as needed.
An expression of the HORWs that exactly solve the NLS equation (1) can be constructed through the following procedure. Consider the expansion coefficients from the suitable series expansion around $\lambda=i$, $F_{\ell}(x,t)$ and $G_{\ell}(x,t),\mathrm{where}\ \ell\in\mathbb{Z}^{+}$, arising from the generating functions
$(1-i\lambda)\frac{\sin\left((x+\lambda t)\sqrt{\lambda^{2}+1}\right)}{\sqrt{\lambda^{2}+1}}$ $=\sum_{\ell=0}^{\infty}\left(\frac{i}{2}\right)^{\ell}F_{\ell}(x,t)(\lambda-i)^{\ell},$
$\cos\left((x+\lambda t)\sqrt{\lambda^{2}+1}\right)$ $=\sum_{\ell=0}^{\infty}\left(\frac{i}{2}\right)^{\ell}G_{\ell}(x,t)(\lambda-i)^{\ell}.$
We use these coefficients to define the following $k\times k$ matrices
$K_{pq}^{(k)}(x,t):=\sum_{\mu=0}^{p-1}\sum_{v=0}^{q-1}\left(\begin{array}[]{c}\mu+v\\ \mu\end{array}\right)\left(F^{*}_{q-v-1}F_{p-\mu-1}+G^{*}_{q-v-1}G_{p-\mu-1}\right),$
and
$H_{pq}^{(k)}(x,t):=-2\left(F_{p-1}+G_{p-1}\right)\left(F^{*}_{q-1}-G^{*}_{q-1}\right),$
where $1\leq p,q\leq k$ and with * denoting complex conjugation, while the arguments of $F$ and $G,$ being $(x,t),$ are implied for brevity.
Rogue waves of order $k$, denoted by $\psi_{k}(x,t),$ are thus furnished by the following formula
$\psi_{k}(x,t):=(-1)^{k}\frac{\operatorname{det}\left(\mathbf{K}^{(k)}(x,t)+\mathbf{H}^{(k)}(x,t)\right)}{\operatorname{det}\left(\mathbf{K}^{(k)}(x,t)\right)}.$ (2)
For example, the $k=1$ Peregrine solution, up to an unimportant phase factor, is given by
$\psi_{1}(x,t):=1-4\frac{1+2\mathrm{i}t}{1+4x^{2}+4t^{2}}.$
The compact representation afforded by Equation (2) coincides with the leading order solution of a parametrized matrix Riemann-Hilbert problem. See  Bilman et al. (2020), in particular Proposition 1, for more details.
We show the space-time evolution of the first four of these rogue waves, in absolute value, in Figure 1. Note that this process can be generalized for arbitrary $k$, leading eventually to the limit of $k\rightarrow\infty$ which corresponds to the waveform discovered in Suleimanov (2017). The sequence of the states, through having $k$ “stems” and from which an extreme focusing emerges, is transparent in the space-time evolution of the patterns in Figure 1.
Figure 1: Visualization of the first four NLS rogue waves, in absolute value, whose formulae are given by Equation (2) and the immediately preceding definitions.
Numerical Examples: the One-Dimensional Case. We now turn to the one-dimensional realization of our proposed experimental protocol (within atomic BECs Pethick and Smith (2002); Pitaevskii and Stringari (2003); Kevrekidis et al. (2015))
We start at the defocusing regime with a parabolic confinement $V(x)=(1/2)\Omega^{2}x^{2}$ and nonlinear defocusing prefactor $g_{\rm defocus}$, for a controllable chemical potential $\mu$; both $\mu$ and $\Omega$ are tunable, and $\mu\gg\Omega$ defines the effective TF regime. The ground state wavefunction is approximately semicircular with density controllably approximated by $|\psi_{TF}|^{2}=\mu-V(x)$, where $\mu>V(x)$. The approximation becomes progressively better as the TF regime is approached, with an accordingly shrinking boundary layer (see, also Karali and Sourdis (2015); Gallo and Pelinovsky (2009)).
Now, at $t=0$, we switch the nonlinearity to a focusing one of prefactor $g_{\rm focus}$, i.e., solving Eq. (1), but with (again, controllable attractive interactions Donley et al. (2001); Roberts et al. (2001); Pollack et al. (2009)) nonlinear prefactor $g_{\rm focus}$ and with this controllably close to a semi-circular initial condition. We use imaginary-time propagation Bao and Du (2003) to compute the ground state with evolution, here and in the focusing case, computed via an adaptize step-size split-step method Sinkin et al. (2003) (see Appendix B for more details). To summarize thhe entire protocol succinctly in two steps:
  • Step 1:
    For a given chemical potential $\mu,$ a trap strength $\Omega,$ and defocusing parameter $g_{\rm defocus}>0$, compute the ground state $\varphi_{0}(x)$ of the defocusing steady state equation $\mu\psi=-\frac{1}{2}{\partial_{x}^{2}\psi}+g_{\rm defocus}|\psi|^{2}\psi+\frac{\Omega^{2}}{2}x^{2}\psi$.
  • Step 2:
    For a given focusing parameter $g_{\rm focus}>0,$ release the trap and use $\varphi_{0}(x)$ as the initial condition in the focusing dynamics given by $i{\partial_{t}\psi}=-\frac{1}{2}{\partial_{x}^{2}\psi}-g_{\rm focus}|\psi|^{2}\psi.$
To make the claim that this protocol enables the controlled generation (as the parameters $\mu$, $g$ and $\Omega$ are varied) of high-order rogue waves, we perform a fit of our numerical realizations thereof, at the time of maximal focusing, to the exact solutions given by Equation (2) rescaled appropriately in space and by the value of the nonlinear focusing coefficient $g_{\rm focus}$ such that the NLS equation (1) remains invariant. We denote the rescaling in space used to minimize the mismatch in an ${L^{2}}$ norm between the numerics and the exact solution by the symbol $s.$ Since the functions given by Equation (2) have a non-zero background, we minimize a mismatch centered at space and with a spatial extent that terminates where the absolute value of $|\psi|$ reaches its last local minimum to the left and to the right of the origin. Denote this spatial region as $\zeta$.
Thus, as a computational problem, the entirety of our work here is a search for the vector $p=\left(\Omega,g_{\rm defocus},g_{\rm focus},\mu,s\right)$ that optimizes (i.e., minimizes) the relative mismatch
$\mathcal{M}(p)=\frac{\left|\|\psi_{p}\|_{L^{2}(\zeta)}^{2}-\|\psi_{k}\|_{L^{2}(\zeta)}^{2}\right|}{\|\psi_{k}\|_{L^{2}(\zeta)}^{2}},$ (3)
where $\psi_{k}$ is the $k^{\rm th}$ order rogue wave given by Equation (2) and $\psi_{p}$ is the wavefunction corresponding to the observed density at the maximal focusing time and generated from our protocol from the 5 dimensional parameter vector $p$. It is important to reiterate here that the wide range of available choices of $\Omega\ll 1$ and $\mu$ are made so as to place us in a suitably elongated 1d regime such that the TF approximation is a meaningful one along this dimension. Such considerations will reflect our choices of trap frequencies and chemical potential also for the higher-dimensional anisotropic traps below). To solve the optimization problem $\min_{p}\mathcal{M}(p),$ we use differential evolution methods, as discussed, e.g., in Storn and Price (1997).
In Figure 2, we report the parameters $p$ found through the minimization and show the match, in absolute-value to facilitate visualization,between the first four generated rogue waves and the exact expressions from Equation (2). Naturally, and similarly to what occurs in the case of the single Peregrine in BECs Romero-Ros et al. (2024), we can only hope to locally match the $k$-th order rogue waves, given the distinct nature of the spatial asymptotics of our initial conditions, hence the need for our proposed matching selection. Nevertheless, our fitting procedure clearly suggests the local spontaneous generation of the relevant multi-hump patterns, by analogy with what happens for Peregrine-based wave breaking in Bertola and Tovbis (2013). Moreover, we demonstrate numerically, in Appendix A, that our experimental protocol shows great promise for remaining robust against uncertainties in the parameter vector $p$ (relevant to potential experimental factors).
Figure 2: Local fitting of our numerical simulations to the first $4$ (i.e., $k=1$ to $k=4$) HORWs given by Equation (2). We report the parameters $p=\left(\Omega,g_{\rm defocus},g_{\rm focus},\mu,s\right)$ to three digits of precision. Top left: $(1.06,0.87,-0.443,3.42,0.914)$. Top right: $(1.23,0.891,-0.701,6.19,1.74)$. Bottom Left: $(1.25,1.07,-0.794,7.42,1.49)$. Bottom Right: $(1.93,1.84,-.483,18.2,1.01)$.
Higher-dimensional Generalizations. Extending considerations to two spatial dimensions, we are able to observe HORWs, e.g., with $k=2$ or $k=3$ in anisotropic variants of the latter setting. More concretely, we show the formation of focusing strongly reminiscent of a second order rogue wave in the top panel of Figure 3. In analogy with our one-dimensional protocol, we start from the defocusing NLS regime of the form:
$i{\partial_{t}\psi}=-\frac{1}{2}{\Delta\psi}+g|\psi|^{2}\psi+V({\bf r})\psi,$ (4)
with $x\in\mathbb{R}^{d}$ and $d=2,3$. It is clear here that we need an experimentally accessible, highly anisotropic 2D trap $V(x,y)=\frac{\Omega_{x}^{2}}{2}x^{2}+\frac{\Omega_{y}^{2}}{2}y^{2}$ in order to achieve the relevant dynamics. We further show the shape of the initial condition, i.e., the ground state of the defocusing problem with $g=g_{\rm defocus}$ used to illustrate the level of anisotropy used to ensure the prevention of finite-time blowup of the dynamics. Similarly to our 1d protocol, at $t=0$, we switch $g=g_{\rm focus}$. During the subsequent focusing dynamics, we release the trap in the $x-$direction while maintaining the trap in the transverse, that is, $y-$direction fixed. In the bottom panel of Figure 3, we show the formation of a $k=3$ rogue wave, as well as an isosurface of the two-dimensional dynamics.
[…]
In three spatial dimensions, anisotropy can once again serve to mitigate the risk of collapse, as is well-known, e.g., from works such as Pérez-García et al. (1997); Salasnich et al. (2002); Márquez and Evaristo (2014). This, in turn, enables us, within the regimes of sufficiently tightly transversely confined BECs discussed therein, to identify parameters that lead to a genuine second order rogue wave-type scenario.
This is shown in Fig. 4 both at the level of the quasi-1D space-time evolution (left panel) and at that of the detailed comparison of the main feature in a quasi-1D integrated form with the 2nd order rogue wave of the Equation (2). In Appendix A, we show that the generated 3D HORW in Fig. 4 is robust under losses as modeled via three body interactions Banerjee et al. (2024).
Conclusions and Future Challenges. In the present work we revisited the framework of higher-order rogue waves that has been recently developed in the mathematical literature (via Darboux transformations and robust inverse scattering methods) as a non-generic, yet well-accessible form of wave breaking. While, to the best of our knowledge, this setup has not been explored at the level of realistic or experimentally relevant settings, we have showcased examples that are fully accessible to current atomic BEC experiments (and also potentially elsewhere, such as, e.g., suitably designed optical media).
We then went on to explore the freedom available to the system through a variation of parameters to provide scenarios enabling the realization of $k=1$ through $k=4$ rogue waves, with a view towards higher-order rogue waves, including the intriguing infinite-order one of Suleimanov (2017). While the relevant idea is, in its essence, tailored to one spatial dimension, we have illustrated that higher-dimensional, highly anisotropic settings are also very much accessible and conducive towards such considerations.
Naturally, this study paves the way for numerous further considerations of related concepts. We only mention a few of these here. One necessity is a wide parametric mapping of the space of possible outcomes of near-semi-circular initial data in the space of $(\mu,\Omega)$ etc. Another question: How amenable are nonintegrable systems to such types of wave breaking? For example, will HORW-type behavior survive in quintic, cubic-quintic, or other recently proposed so-called droplet Luo et al. (2021) variants of the model? A further direction stems from spatially discrete models and concerns whether a variant of such higher-order (including infinite-order Suleimanov (2017)) rogue waves exists of Ablowitz-Ladik type and discrete nonlinear Schrödinger type lattices Ablowitz et al. (2004); Kevrekidis (2009).
It can thus be inferred that this theme presents a wide palette of opportunities in one- and higher- spatial dimensions that are extremely timely and relevant for near-future analytical, numerical and experimental considerations.
Acknowledgements. This material is based upon work supported by the U.S. National Science Foundation under the awards PHY-2110030, PHY-2408988, DMS-2204702 (PGK), and PHY-2316622 (JA). The authors gratefully acknowledge the Society for Industrial and Applied Mathematics (SIAM) postdoctoral support program, established by Martin Golubitsky and Barbara Keyfitz, for helping make this work possible. PGK is also grateful to Professor Simos Mistakidis for relevant discussions.
Appendix A Robustness Studies
Here we report two robustness studies of the numerically simulated one-dimensional HORWs reported in the main text. The first study we conduct is as follows. We randomly perturb each component of the optimal parameter vector $p^{*},$ for the parameters reported in the caption of Figure 2, by at most 1% of each parameter value using uniform sampling. In Figure 5, we show a histogram plot for the realized mismatches $\mathcal{M}(p^{*},w_{p})$, where $w_{p}$ is the realized noise such that each component of the vector $p+w_{p}$ is within 1% of their respective values. We then repeat the same study for an uncertainty of 10%, omitting the histogram in this case for brevity. We report the histogram at 1% to show that the numerical realizations shown in Figure 6 are not at the extreme of the left tail of the empirical distribtuion of HORW realizations.
The main conclusion that we draw here is that, especially for lower order rogue waves, small perturbations do not significantly shift the mismatch, i.e., nearby parameters yield similarly favorable mismatches. However, even when the mismatches tend to be large, especially at the level of 10% uncertainty, we still observe that the HORWs retain their overall shape. Our numerics predict that these generated HORWs would remain significantly detectable in an experimental setting despite the relatively large uncertainty in the tuning parameters of the scheme. In Figure 6, we show a few realizations of the resulting 1% perturbed HORW against their zero noise counterparts from Figure 2. We also show the resulting 10% perturbed HORW against their zero noise counterparts in Figure 7.
Figure 5: Histogram plots of the realized mismatches for perturbed parameters $p^{*}+w_{p}$ (with a 1% uncertainty) corresponding with Figure 2 in the main text. Top Left: $k=1$. Top Right: $k=2$. Bottom Left: $k=3$. Bottom Right: $k=4$.
Figure 6: Three realizations, in absolute-value, of HORWs, sequentially in order, generated from $p^{*}+w_{p}$ at 1% of uncertaintiy, and corresponding with Figure 2 in the main text. Top Left: $k=1$. Top Right: $k=2$. Bottom Left: $k=3$. Bottom Right: $k=4$.
Figure 7: Three realizations, in absolute-value, of HORWs, sequentially in order, generated from $p^{*}+w_{p}$ at 10% of uncertaintiy, and corresponding with Figure 2 in the main text. Top Left: $k=1$. Top Right: $k=2$. Bottom Left: $k=3$. Bottom Right: $k=4$. Despite the larger changes in the mismatch, as compared with Figure 7, the excited rogue waves maintain their overall shape.
The second study we conduct considers three-body losses. We model these effects when the focusing dynamics take place, as this is where the losses become most relevant due to the extreme focusing. The focusing dynamics we consider are given by
$i{\partial_{t}\psi}=-\frac{1}{2}{\Delta\psi}-g_{\rm focus}|\psi|^{2}\psi-iK_{3}|\psi|^{4}\psi.$ (5)
In Figure 8 we show the functional dependence of the mismatch on the parameter $K_{3}$. In Figure 9, we show examples of how the losses impact the focusing of the dynamics into each HORW in one dimension. We also show the impact of the losses in a three dimensional scenario in Figure 10.
It should be kept in mind when considering these figures that, using realistic numbers such as those of Ref. Banerjee et al. (2024), the dimensionless value of $K_{3}$ turns out to be quite low, roughly about $10^{-3}-10^{-4}$ for the setup reported in Banerjee et al. (2024). Hence, we expect the modification of the reported mismatch and the relevant profile variation to be rather minimal based on the data of Figs. 8 and 9. In general, however, it should be also observed that even for much more substantial values of the dissipation, the relevant graphs of 9 and 10 are qualitatively quite similar to the ones without the damping effects and hence our phenomenology is fairly robust under the presence of realistic three-body losses.
Figure 8: Functional dependence of the mismatch on the parameter $K_{3}$ from losses as modeled by Equation (5). The optimal parameter vector $p^{*}$ is the one reported in Figure 2. Top Left: $k=1$. Top Right: $k=2$. Bottom Left: $k=3$. Bottom Right: $k=4$.
Figure 9: Realized HORWs for different values of $K_{3}$ from losses as modeled by Equation (5). The zero loss cases are the ones reported in Figure 2. Top Left: $k=1$. Top Right: $k=2$. Bottom Left: $k=3$. Bottom Right: $k=4$.
Figure 10: Realized second order rogue wave with a loss modeled by Equation (5). The parameter $K_{3}=0.005$ shows that a relatively small loss does not distort the shape of the generated rogue wave drastically. The isocontour on the right was set to 12, consistent with the zero loss case shown in Figure 4.
Appendix B GPE Simulation Technique
To simulate the lossy GPE (5), we use a standard operator splitting technique. That is, we rewrite Equation (5) as
$i\partial_{t}=\mathcal{L}\psi+\mathcal{N}(\psi)$
where
$\mathcal{L}\psi=-\frac{1}{2}{\Delta\psi},\qquad\mathcal{N}(\psi)=-g|\psi|^{2}\psi-iK_{3}|\psi|^{4}\psi.$
We diagonalize the matrix exponential of the linear operator numerically via fast Fourier/inverse Fourier transforms.
The nonlinearity $\mathcal{N}(\psi)$ can be handled using basic quadrature techniques. First, we use polar coordinates, that is,
$\psi=\rho e^{i\theta}$
The resulting equations for the phase and amplitude are
$\dot{\theta}=g\rho^{2},\qquad\dot{\rho}=-K_{3}\rho^{5}$
Integrating the equation for the amplitude from $t_{n}$ to $t_{n+1}$ is done exactly;
$\rho_{n+1}=\frac{\rho_{n}}{\sqrt[4]{1+4K_{3}h_{n}\rho_{n}}},$
where $h_{n}$ is the stepsize $t_{n+1}-t_{n}$. Integrating the equation for the phase can be handled using the trapezoidal rule:
[…]
Simplifying using the fact that $\rho_{n}=|\psi_{n}|$ and $\theta_{n}=\mathrm{Arg}\ \psi_{n}$, we have the locally third-order in time update
$\psi_{n+1}=\rho_{n+1}e^{i\theta_{n+1}}$
where
$\rho_{n+1}=\frac{|\psi_{n}|}{\sqrt[4]{1+4K_{3}h_{n}|\psi_{n}|}}$
$\theta_{n+1}=\mathrm{Arg}\ \psi_{n}+\frac{gh_{n}}{2}|\psi_{n}|^{2}\left(1+(1+4K_{3}h_{n}|\psi_{n}|)^{-1/2}\right)$
Note that when $K_{3}=0$, the nonlinear update reduces to
$\rho_{n+1}=|\psi_{n}|,\qquad\theta_{n+1}=\mathrm{Arg}\ \psi_{n}+gh_{n}|\psi_{n}|^{2},$
and is exact within the evaluation of the matrix operator of the nonlinearity. This scheme is what is used in the numerical solution of the focusing and trap free $(V\equiv 0)$ version of Equation (4).
To make the method adaptive and globally third-order in time for the efficient resolution of the transient and extreme focusing that we observe throughout this work, we use a combination of Strang Splitting and Richardson extrapolation; see Sinkin et al. (2003) for an excellent explanation of this method. We use a similar strategy to compute the ground state of the defocusing equation using a renormalized imaginary time propagation method; please see Bao and Du (2003) for more details.